A stock sells for $30. The next dividend will be $6 per share. If the return on equity ROE is a constant 15% and the company reinvests 20% of earnings in the firm, what must be the opportunity cost of capital? (Do not round intermediate calculations.)

Cost of Capital %
Answer #1

b = retention ratio = 20%, ROE = 15%

Growth rate = g = b x ROE = 20% x 15% = 3%

ke = D1/P0 + g = 6/30 + 0.03 = 0.20 + 0.03 = 0.23 = 23%

Answer : 23% (Thumbs up please)